Having Trouble understanding the procedure for finding the answer.

1) Solve the system using substitution of linear equations.
the 2 linear equations are:
a)2x + 3y = -4 b)y - 2x = 4

thanks

Answer by stanbon(75887)

Solve the system using substitution of linear equations.
the 2 linear equations are:
a)2x + 3y = -4
b)y - 2x = 4
-------------------
y = 2x+4
---
Substitute into the other equation and solve for "x":
2x + 3(2x+4) = -4
8x + 12 = -4
8x = -16
x = -2
------
Since y = 2x+4, y = 2(-2)+4 = 0
-----------------------------------
Solution (-2,0)
====================
Cheers,
Stan H.